# Secret Breach Detection in Source Code with Large Language Models

### v1

@@ II-C3 Evaluation Data
TABLE X: Evaluation Metrics for Zero-Shot Prompting (Fine-Tuned Models)
Model Precision Recall F1-Score
DeepSeek-7B 0.9503 0.9473 0.9472
Gemma-7B 0.9756 0.9753 0.9753
LLaMA-3.1-8B 0.9861 0.9853 0.9852
Mistral-7B 0.9815 0.9825 0.9819
CodeLLaMA-7B 0.9716 0.9709 0.9707
II-C4 Result
@@
The performance of all five LLMs—DeepSeek, Gemma, LLaMA-3.1-8B, Mistral, and CodeLLaMA—under different prompting settings without fine-tuning is summarized in Table VII, Table VIII, and Table IX. Table X compares their performance in the zero-shot setting after fine-tuning. Across all scenarios, the results clearly show that fine-tuning significantly boosts performance. Among the pre-trained models, LLaMA-3.1 8B consistently achieves the best F1-scores in all prompting conditions.
@@
  • Zero-Shot (Raw): As seen in Table VII, model performance is relatively low, with F1-scores ranging from 0.33 to 0.52. CodeLLaMA-7B performs the best in this setting with an F1-score of 0.5207.
@@
  • One-Shot (Raw): Table VIII shows a noticeable improvement in performance. LLaMA-3.1-8B achieves the highest F1-score of 0.6699, followed by Gemma and CodeLLaMA.
@@
  • Few-Shot (Raw): With more contextual examples, performance improves further (Table IX). LLaMA-3.1 8B again leads with an F1-score of 0.7913, followed by Mistral-7B and Gemma-7B.
@@
  • Zero-Shot (Fine-Tuned): As shown in Table X, fine-tuning significantly enhances model effectiveness. LLaMA-3.1-8B attains the top F1-score of 0.9852, followed by Gemma and CodeLLaMA. Even DeepSeek-7B, which had low performance in the raw setting, achieves an impressive 0.9472 F1-score after fine-tuning.
@@
The confusion matrix in Figure 4 illustrates the performance of the fine-tuned LLaMA-3.1-8B model under the zero-shot prompting in fine-tuned model. The model correctly identified 1,496 out of 1,500 negative (non-secret) instances and 1,463 out of 1,500 positive (secret) instances. Only 4 false positives and 37 false negatives were observed, indicating high precision and recall. These results are consistent with the high F1-score (0.9852) observed in Table X, affirming the model’s strong discriminative ability after fine-tuning. The near-diagonal dominance of the matrix further emphasizes the model’s robustness in accurately separating secret and non-secret examples.
@@
II-C5 Multi class classification
To overcome the limitations of pretrained LLMs, we fine-tuned smaller language models on 9000 labeled positive samples, specifically curated for the secret classification task. These models were then evaluated on the same 3000-sample test set used for the pretrained models to ensure a fair comparison.
TABLE XI: Overall Multiclass Classification Accuracy
Model F1-Score(Weighted)
Deepseek-7B 0.9584
Gemma-7B 0.9494
Codellama-7B 0.9625
LLaMA-3.1-8B 0.980
Mistral-7B 0.982
TABLE XII: Multiclass Classification Report for Mistral-7B
Category Precision Recall F1-Score
Private Key 0.9875 0.9893 0.9884
API Key and Secret 0.9879 0.9760 0.9819
Authentication Key and Token 0.9972 0.9945 0.9959
Other 0.8788 0.8878 0.8832
Generic Secret 0.8060 0.9153 0.8571
Database and Server URL 1.0000 1.0000 1.0000
Password 0.9512 1.0000 0.9750
Username 1.0000 1.0000 1.0000
Overall Accuracy 0.9820
Table XII presents the classification reports for Mistral model. Figure 5 presents the confusion matrix for the Mistral-7B model, providing insights into its classification performance across different secret types. The model performs well on high-frequency categories, such as Private Key (1109 correct), API Key and Secret (896 correct), and Authentication Key and Token (721 correct), indicating a strong ability to distinguish these secret types.
Figure 5: Confusion Matrix for Mistral-7B fine tuned for multi class classification
While overall accuracy and weighted f1-scores are high, some misclassifications are observed. API Key and Secret secrets are occasionally predicted as Other or Authentication Key and Token, suggesting feature overlap. Similarly, the Other category exhibits greater misclassifications, reflecting its diverse nature. Lower-frequency categories, such as Generic Secret, Database and Server URL, and Password, are correctly classified in most cases, but minor errors remain.
RQ2 Summary: We evaluated the effectiveness of fine-tuned open-source LLMs (e.g., LLaMA-3.1 8B, Mistral, CodeLLaMA, DeepSeek, and Gemma) for secret detection using both binary and multiclass classification. Fine-tuning significantly improved performance, with LLaMA-3.1 8B achieving the highest F1 score (0.9852) in zero-shot settings. In multiclass classification, both LLaMA-3.1 8B and Mistral-7B achieved high accuracy (98%+), with Mistral showing strong category-wise precision and recall. These results affirm that compact, fine-tuned models are highly effective and resource-efficient alternatives to commercial LLMs for sensitive software engineering tasks like secret detection.
III Discussion
III-A Analysis of Detected Secrets
@@
III-B Comparison with Existing Tools
@@
The best-performing fine-tuned LLaMA-3.1 8B model achieved an F1-score of 0.9852 on our subset of 3,000 samples (1,500 secrets and 1,500 non-secrets). In contrast, a regex-only tool would flag all candidates as secrets, yielding a theoretical precision of just 50% in this balanced setting.
@@ III-C False Negatives
Despite the strong performance of our fine-tuned models, a limited number of false negatives were observed during evaluation. Specifically, 37 instances were incorrectly classified as non-sensitive. A detailed manual analysis revealed several contributing factors behind these misclassifications.
@@
One source of error was inconsistencies between our annotation framework and the labeling conventions of the SecretBench dataset [24]. For example, the dataset labels certain public keys—such as consensus public key—as sensitive, while our model and prompt design treated such keys as non-sensitive based on commonly accepted definitions of sensitive information. As a result, some predictions were considered false negatives relative to the dataset, although they were consistent with our defined classification criteria. A similar discrepancy was found in cases such as pk_test_..., which our model interpreted as a benign public key based on its naming pattern and prior examples. However, the dataset labeled it as sensitive, potentially reflecting real-world risks associated with leaked test tokens, even when such keys are publicly documented to be accessible from the client side (e.g., Stripe documentation). These examples highlight the ambiguity that can exist in the classification of security-relevant identifiers.
Another contributing factor was the model’s limited capacity to accurately detect secrets when the surrounding context was insufficient. This limitation was particularly evident in cases involving long-format secrets, such as RSA private keys, which often exceed the default input length. To mitigate this, we increased the model’s maximum input sequence length from 512 to 1024 tokens. This modification resulted in a measurable improvement, reducing the number of false negatives by enabling the model to process a more complete context for each candidate.
@@
Additionally, we observed that very small number of false negatives arose due to subtle variations in context. Identical candidate strings appeared in different code environments and received inconsistent classifications. This behavior suggests that the model may be sensitive to minor contextual cues, leading to inconsistent decision-making. This represents a potential area for future improvement, particularly in enhancing the model’s robustness to contextual variability.
Overall, we acknowledge that the presence of a small number of false negatives reflects a trade-off aimed at significantly reducing false positives. By aligning the model’s predictions with a clearly defined set of criteria, we prioritized precision in sensitive data detection. In applications where minimizing false alarms is essential to maintaining developer trust and usability, we believe this trade-off is both justified and desirable.
III-D Impact of Imbalanced Training Distribution
@@
In addition to balanced training, we also fine-tuned our models using an imbalanced training set that better reflects the real-world distribution of secrets in source code. Specifically, we used a training dataset comprising 3,750 true secrets and 20,250 non-secret samples, maintaining the total size at 24,000 instances. This distribution mirrors the actual ratio observed in the SecretBench dataset, where only 15,084 out of 97,479 samples are true secrets.
@@
Despite the skewed class proportions, the performance of the fine-tuned models remained robust. As shown in Table XIII, our best-performing model, LLaMA-3.1-8B, achieved a precision of 0.9742, recall of 0.9730, and an F1-score of 0.9730 on a balanced evaluation set. When trained on the imbalanced dataset, these scores experienced only a slight drop (1–3% across most metrics), indicating that LLMs retain their contextual reasoning capabilities and strong classification performance even under class imbalance.
This resilience can be attributed to the expressive power of LLMs in understanding code semantics and distinguishing true secrets from non-sensitive tokens based on context. Unlike traditional classifiers that may suffer significantly in imbalanced scenarios, LLMs leverage large-scale pretraining and fine-grained token-level understanding to preserve high recall without compromising much on precision.
Our findings confirm that even when trained on datasets resembling real-world secret distributions, fine-tuned LLMs can maintain excellent detection capabilities. This makes them especially suitable for deployment in practical settings, where secrets are rare but their detection is critical.
@@
TABLE XIII: Evaluation Metrics for Models Fine tuned on Imbalanced dataset
Model Precision Recall F1-Score
DeepSeek-7B 0.9241 0.9117 0.9110
Gemma-7B 0.9656 0.9451 0.9523
LLaMA-3.1-8B 0.9742 0.9730 0.9730
Mistral-7B 0.9715 0.9724 0.9720
CodeLLaMA-7B 0.9571 0.9533 0.9532
III-E Analysis of Impact of Context Window
@@
Table XIV shows the F1-scores obtained by each fine-tuned model when tested under these two context window sizes.
TABLE XIV: Impact of Context Window Size on F1-Score (Fine-Tuned Models)
Model F1-Score (200 chars) F1-Score (300 chars)
DeepSeek-7B 0.9472 0.9690
Gemma-7B 0.9753 0.9764
LLaMA-3.1-8B 0.9852 0.9882
Mistral-7B 0.9819 0.9861
CodeLLaMA-7B 0.9707 0.9721
The results indicate that increasing the context window from 200 to 300 characters leads to a modest improvement in F1-score across all models. This improvement is attributed to the additional contextual cues available to the model for making more informed classifications—especially useful for complex or multiline secrets.
@@
III-F Inference Time Analysis
We evaluated inference speed using 3,000 sequential predictions on a local GPU setup. The average inference time per sample was 0.1924 seconds on an RTX 4090, demonstrating that our models support real-time or near-real-time deployment in environments like IDEs or CI/CD pipelines.
III-G Resource Efficiency and Cost
All experiments ran on a workstation with an Intel i5-13400F CPU, 128 GB RAM, and an RTX 4090 GPU. Thanks to QLoRA and PEFT, we fine-tuned 7B–8B models efficiently within local hardware limits. This enables high-performing secret detection without costly cloud infrastructure, supporting privacy-focused and budget-conscious deployments.
IV Related Work
Large-Scale Secret Leak Analysis
Meli et al. [15] conducted a large-scale study on secret leaks within GitHub by analyzing billions of files through real-time scans and snapshots. They found that over 100,000 repositories expose thousands of secrets daily. Their regex, entropy-based filtering, and pattern detection approach achieved 99.29% accuracy.
Program Analysis for Secret Detection
Building upon earlier approaches, Sinha et al. [16] enhanced secret detection by incorporating program slicing, keyword-based, and pattern-based searches. Their method achieved high precision and recall for detecting API keys while also addressing false positives and the limitations of proximity-based detection.
Machine Learning-Based Filtering
To further reduce false positives, Saha et al. [17] introduced machine learning-based methods. They created a dataset of secrets and non-secrets from GitHub using regex and trained models on binary features like character patterns and entropy. Their Soft Voting Classifier achieved an F1-score of 86.7%.
Deep Learning Approaches
Feng et al. [19] explored deep learning techniques for detecting secrets, particularly textual passwords. By leveraging deep neural networks, they captured semantic patterns and significantly reduced false positives, demonstrating deep learning’s advantages over traditional rule-based methods.
Secret Detection in Issue Reports
Secret Mitigation Techniques
In terms of mitigation, Sinha et al. [40] suggested strategies like excluding secrets from version control, encrypting them using tools such as git-crypt, and integrating automated scanning mechanisms. They also emphasized practices like commit signing and two-factor authentication (2FA) to strengthen security.
Best Practices for Secret Management
Best practices for secret management were compiled by Basak et al. [23], who proposed 24 actionable guidelines. These included avoiding hard-coded secrets, using environment variables, employing short-lived credentials, utilizing secret managers like Vault, and sanitizing version control system (VCS) history.
Comparative Analysis of Detection Tools
V Threats to Validity

### v2

@@ IV-B Secret Detection Using Small Large Language Models (RQ2)
TABLE V: Evaluation Metrics for Prompting Settings on Raw and Fine-Tuned Models (Zero-, One-, Few-Shot)
Setting Model Precision Recall F1-Score F2-Score
Zero- Shot (Raw) DeepSeek-7B 0.4105 0.4283 0.4054 0.4246
Gemma-7B 0.5501 0.5003 0.3353 0.5095
LLaMA-3.1-8B 0.5412 0.5130 0.4125 0.5184
Mistral-7B 0.5033 0.5033 0.5029 0.5033
CodeLLaMA-7B 0.5453 0.5383 0.5207 0.5397
One- Shot (Raw) DeepSeek-7B 0.6864 0.5743 0.4990 0.5937
Gemma-7B 0.6741 0.6610 0.6545 0.6636
LLaMA-3.1-8B 0.7622 0.6920 0.6699 0.705
Mistral-7B 0.6848 0.5720 0.4949 0.5915
CodeLLaMA-7B 0.6741 0.6610 0.6545 0.6636
Few- Shot (Raw) DeepSeek-7B 0.5216 0.5127 0.4563 0.5145
Gemma-7B 0.7286 0.7187 0.7156 0.7207
LLaMA-3.1-8B 0.7913 0.7913 0.7913 0.7913
Mistral-7B 0.7678 0.7630 0.7619 0.764
CodeLLaMA-7B 0.7286 0.7187 0.7156 0.7207
Zero- Shot (Fine- tuned) DeepSeek-7B 0.9503 0.9473 0.9472 0.9479
Gemma-7B 0.9756 0.9753 0.9753 0.9754
LLaMA-3.1-8B 0.9861 0.9853 0.9852 0.9855
Mistral-7B 0.9815 0.9825 0.9819 0.9823
CodeLLaMA-7B 0.9716 0.9709 0.9707 0.971
@@
The performance of all five LLMs—DeepSeek, Gemma, LLaMA-3.1-8B, Mistral, and CodeLLaMA—under different prompting settings without fine-tuning and after fine-tuning is summarized in Table V. Across all scenarios, the results clearly show that fine-tuning significantly boosts performance. Among the pre-trained models, LLaMA-3.1 8B consistently achieves the best F1-scores in all prompting conditions.
@@
  • Zero-Shot (Raw): As seen in Table V, model performance is relatively low, with F1-scores ranging from 0.33 to 0.52. CodeLLaMA-7B performs the best in this setting with an F1-score of 0.5207 and F2-score of 0.5397.
@@
  • One-Shot (Raw): LLaMA-3.1-8B achieves the highest F1-score of 0.6699 and F2-score 0.705, followed by Gemma and CodeLLaMA.
@@
  • Few-Shot (Raw): With more contextual examples, performance improves further. LLaMA-3.1-8B again leads with F1 and F2-scores of 0.7913, followed by Mistral-7B and Gemma-7B.
@@
  • Zero-Shot (Fine-Tuned): Fine-tuning significantly enhances model effectiveness. LLaMA-3.1-8B attains the top F1-score of 0.9852 and F2-score of 0.9855, followed by Gemma and CodeLLaMA. Even DeepSeek-7B, which had low performance in the raw setting, achieves an impressive 0.9472 F1-score after fine-tuning.
@@
The confusion matrix in Figure 4 illustrates the performance of the fine-tuned LLaMA-3.1-8B model under the zero-shot prompting in fine-tuned model. The model correctly identified 1,496 out of 1,500 negative (non-secret) instances and 1,463 out of 1,500 positive (secret) instances. Only 4 false positives and 37 false negatives were observed, indicating high precision and recall. These results are consistent with the high F1-score (0.9852) observed in Table V, affirming the model’s strong discriminative ability after fine-tuning. The near-diagonal dominance of the matrix further emphasizes the model’s robustness in accurately separating secret and non-secret examples.
@@
IV-B1 Multiclass Classification Performance
Table VI presents the overall weighted F1-scores for various LLMs fine-tuned for multiclass classification. Among the evaluated models, Mistral-7B achieves the highest weighted F1-score of 0.982, closely followed by LLaMA-3.1-8B and Codellama-7B.
TABLE VI: Overall Multiclass Classification Accuracy
Model F1-Score (Weighted) F2-Score(Weighted)
Deepseek-7B 0.9584 0.9581
Gemma-7B 0.9494 0.9472
CodeLLama-7B 0.9625 0.9618
LLaMA-3.1-8B 0.9800 0.9798
Mistral-7B 0.9820 0.9822
To further analyze the best-performing model, Table VII presents the detailed classification report for Mistral-7B across all secret types. Figure 5 shows its confusion matrix, highlighting per-category prediction strengths and weaknesses.
TABLE VII: Multiclass Classification Report for Mistral-7B
Category Precision Recall F1-Score F2-Score
Private Key 0.9875 0.9893 0.9884 0.9889
API Key and Secret 0.9879 0.9760 0.9819 0.9784
Authentication Key and Token 0.9972 0.9945 0.9959 0.9950
Other 0.8788 0.8878 0.8832 0.8860
Generic Secret 0.8060 0.9153 0.8571 0.8911
Database and Server URL 1.0000 1.0000 1.0000 1.0000
Password 0.9512 1.0000 0.9750 0.9898
Username 1.0000 1.0000 1.0000 1.0000
Weighted F1-score 0.9820
Weighted F2-score 0.9822
Figure 5: Confusion Matrix for Mistral-7B fine-tuned for multiclass classification
The Mistral-7B model excels in high-frequency categories such as Private Key (1,109 correct), API Key and Secret (896), and Authentication Key and Token (721), showing strong discriminative ability. Some misclassifications occur, particularly between API Key and Secret, Other, and Authentication Key and Token, likely due to feature overlap. Even lower-frequency categories like Generic Secret, Database and Server URL, and Password are accurately classified in most cases.
RQ2 Summary: We evaluated the effectiveness of fine-tuned open-source LLMs (e.g., LLaMA-3.1-8B, Mistral, CodeLLaMA, DeepSeek, and Gemma) for secret detection using both binary and multiclass classification. Fine-tuning significantly improved performance, with LLaMA-3.1-8B achieving the highest F1 and F2-score (0.9852 and 0.9855) in zero-shot settings. In multiclass classification, both LLaMA-3.1-8B and Mistral-7B achieved high F1 and F2-scores (98%+), with Mistral showing strong category-wise precision and recall. These results affirm that compact, fine-tuned models are highly effective and resource-efficient alternatives to commercial LLMs for sensitive software engineering tasks like secret detection.
V Discussion
V-A Analysis of Detected Secrets
@@
V-B Comparison with Existing Tools
@@
The best-performing fine-tuned LLaMA-3.1-8B model achieved an F1-score of 0.9852 on our subset of 3,000 samples (1,500 secrets and 1,500 non-secrets). In contrast, a regex-only tool would flag all candidates as secrets, yielding a theoretical precision of just 50% in this balanced setting.
@@ V-C False Negatives
Despite strong performance, our fine-tuned models produced 37 false negatives during evaluation. Manual analysis revealed key causes behind these misclassifications.
@@
One issue stemmed from inconsistencies between our annotation framework and SecretBench’s labeling conventions [23]. For example, the dataset labels certain public keys like consensus public key as sensitive, whereas our model treated them as non-sensitive based on our defined classification criteria.. A similar discrepancy was found in cases such as pk_test_..., which our model interpreted as a benign public key based on its naming pattern and prior examples. However, the dataset labeled it as sensitive, potentially reflecting real-world risks associated with leaked test tokens, even when such keys are publicly documented to be accessible from the client side (e.g., Stripe documentation). These cases highlight ambiguity in defining sensitive identifiers.
Another factor was the model’s limited ability to detect secrets with insufficient context—particularly for long-format secrets like RSA keys, which exceeded the default input size. Increasing the input length from 512 to 1024 tokens improved performance by reducing such false negatives.
@@
A very small number of false negatives arose due to subtle variations in context. Identical candidate strings appeared in different code environments and received inconsistent classifications. This suggests that the model may be sensitive to minor contextual cues, leading to inconsistent decision-making. This represents a potential area for future improvement, particularly in enhancing the model’s robustness to contextual variability.
Overall, the limited false negatives reflect a trade-off favoring lower false positives. By aligning with strict classification criteria, we prioritized precision, which is crucial for maintaining developer trust and usability in practical deployments.
V-D Impact of Imbalanced Training Distribution
@@
Alongside balanced training, we also fine-tuned our models using an imbalanced dataset that mirrors the real-world distribution of secrets in source code. This dataset included 3,750 true secrets and 20,250 non-secret samples, totaling 24,000 instances—closely reflecting the SecretBench ratio, where only 15,084 out of 97,479 samples are actual secrets.
@@
Despite this class imbalance, our models maintained strong performance. As shown in Table VIII, LLaMA-3.1-8B achieved a precision of 0.9742, recall of 0.9730, F1-score of 0.9730 and F2-score of 0.9732 when evaluated on a balanced test set. When trained on the imbalanced dataset, these metrics declined only slightly (by 1–3%), indicating that LLMs can generalize well even with skewed training distributions.
This robustness can be attributed to the contextual understanding and token-level reasoning inherent in LLMs, which allows them to distinguish secrets from benign strings based on nuanced patterns. Unlike traditional classifiers that often degrade under imbalance, LLMs retain both high recall and precision thanks to their extensive pretraining and semantic awareness.
These findings suggest that LLM-based models remain highly effective even when trained under realistic, imbalanced conditions. This makes them particularly suitable for real-world deployments, where secrets are rare but detecting them accurately is critical for maintaining software security.
@@
TABLE VIII: Evaluation Metrics for Models Fine-tuned on Imbalanced Dataset
Model Precision Recall F1-Score F2-Score
DeepSeek-7B 0.9241 0.9117 0.911 0.9142
Gemma-7B 0.9656 0.9451 0.9523 0.9491
LLaMA-3.1-8B 0.9742 0.973 0.9730 0.9732
Mistral-7B 0.9715 0.9724 0.9720 0.9722
CodeLLaMA-7B 0.9571 0.9533 0.9532 0.9541
V-E Analysis of Impact of Context Window
@@
Table IX shows the F1-scores obtained by each fine-tuned model when tested under these two context window sizes.
TABLE IX: Impact of Context Window Size on F1 and F2 Scores (Fine-Tuned Models)
Model F1 (200) F2 (200) F1 (300) F2 (300)
DeepSeek-7B 0.9472 0.9601 0.9690 0.9784
Gemma-7B 0.9753 0.9809 0.9764 0.9815
LLaMA-3.1-8B 0.9852 0.9895 0.9882 0.9887
Mistral-7B 0.9819 0.9870 0.9861 0.9883
CodeLLaMA-7B 0.9707 0.9755 0.9721 0.9767
The results indicate that increasing the context window from 200 to 300 characters leads to a modest improvement in F1 and F2-score across all models. This improvement is attributed to the additional contextual cues available to the model for making more informed classifications—especially useful for complex or multiline secrets.
@@
V-F Training Setup
All fine-tuning experiments were conducted on a local workstation equipped with an Intel i5-13400F CPU, 128 GB RAM, and an NVIDIA RTX 4090 GPU. We utilized QLoRA and PEFT techniques to enable memory-efficient training of 7B–8B models. This setup ensures that the experiments can be reproduced on advanced consumer-grade hardware without requiring access to large-scale cloud infrastructure.
V-G Training and Inference Efficiency
We evaluated the resource usage and performance of each model during both training and inference. Despite variations in model size and architecture, GPU VRAM usage during training remained below 17GB. As shown in Table X, training time ranged from 16 to 25 hours, and inference memory usage stayed below 13GB. Inference latency per sample varied slightly across models, with the fastest being DeepSeek-7B at 0.091 seconds. These results highlight the feasibility of integrating these fine-tuned models into real-time or near-real-time developer environments.
TABLE X: Resource Efficiency Metrics for Fine-Tuned Models
Model Train VRAM (GB) Train Time (hrs) Inference VRAM (GB) Inference Time (s)
DeepSeek-7B 9.95 16.3 10.18 0.091
Gemma-7B 16.3 23.5 12.02 0.112
LLaMA-3.1-8B 15.02 24.3 9.08 0.100
Mistral-7B 10.08 24.75 8.21 0.136
CodeLLaMA-7B 9.98 22.5 10.35 0.122
VI Threats to Validity